# Quantifying the Limits of Segmentation Foundation Models: Modeling Challenges in Segmenting Tree-Like and Low-Contrast Objects

### v2

@@ 4.3.2 Experiments on Real Data
DIS antenna branch fence hanger log wire
Model $\tau$ $\rho$ $\tau$ $\rho$ $\tau$ $\rho$ $\tau$ $\rho$ $\tau$ $\rho$ $\tau$ $\rho$ $\tau$ $\rho$
CPR: SAM, ViT-H [26] -0.59 -0.76 -0.13 -0.21 -0.61 -0.81 -0.61 -0.81 -0.25 -0.37 -0.60 -0.77 -0.06 -0.10
SAM, ViT-B [26] -0.63 -0.81 -0.18 -0.27 -0.63 -0.82 -0.63 -0.83 -0.44 -0.62 -0.62 -0.80 -0.12 -0.17
SAM 2, ViT-L [36] -0.41 -0.57 -0.30 -0.43 -0.64 -0.84 -0.61 -0.81 -0.48 -0.66 -0.67 -0.85 -0.14 -0.21
SAM 2, ViT-B+ [36] -0.38 -0.52 -0.16 -0.23 -0.52 -0.71 -0.57 -0.78 -0.42 -0.59 -0.60 -0.77 -0.12 -0.18
HQ-SAM, ViT-H [24] -0.52 -0.69 -0.23 -0.33 -0.77 -0.93 -0.79 -0.95 -0.51 -0.70 -0.69 -0.85 -0.22 -0.33
HQ-SAM, ViT-B [24] -0.62 -0.81 -0.25 -0.34 -0.75 -0.92 -0.83 -0.96 -0.50 -0.69 -0.71 -0.87 -0.24 -0.35
Average -0.53 -0.69 -0.21 -0.30 -0.65 -0.84 -0.67 -0.86 -0.43 -0.61 -0.65 -0.82 -0.15 -0.22
DoGD: SAM, ViT-H [26] 0.58 0.77 0.23 0.30 0.52 0.70 0.46 0.65 0.52 0.72 0.30 0.46 0.08 0.11
SAM, ViT-B [26] 0.45 0.64 0.12 0.14 0.51 0.69 0.45 0.64 0.50 0.70 0.13 0.22 0.06 0.08
SAM 2, ViT-L [36] 0.44 0.60 0.42 0.59 0.48 0.65 0.49 0.68 -0.05 -0.08 0.56 0.75 -0.03 -0.05
SAM 2, ViT-B+ [36] 0.26 0.39 0.40 0.56 0.29 0.40 0.30 0.44 -0.06 -0.10 0.42 0.58 -0.08 -0.12
HQ-SAM, ViT-H [24] 0.48 0.66 0.53 0.72 0.58 0.72 0.48 0.67 -0.31 -0.45 0.57 0.75 -0.13 -0.19
HQ-SAM, ViT-B [24] 0.50 0.69 0.67 0.85 0.59 0.73 0.57 0.77 -0.25 -0.34 0.53 0.71 -0.04 -0.06
Average 0.45 0.63 0.40 0.53 0.50 0.65 0.46 0.64 0.06 0.08 0.42 0.58 -0.02 -0.04
Table 2: Rank correlation between SFM IoU and object tree-likeness (CPR, upper table and DoGD, lower table), on DIS and iShape.
Figure 5: Segmentation IoU vs. object tree-likeness (CPR, top; DoGD, bottom) on DIS (left) for all three SFMs and iShape (right) shown for SAM.
We see that despite the many potential confounding factors in real data, there is still a clear correlation between object tree-likeness as measured by the proposed metrics and SFM segmentation performance. In particular, we see average absolute correlations of $|\tau|=0.59$ and $|\rho|=0.76$ for CPR and $|\tau|=0.38$ and $|\rho|=0.51$ for DoGD—excluding the antenna and wire objects of iShape which had outlying correlations likely because those two object classes cover only small ranges of tree-likeness as opposed to the other iShape classes (Fig. 5), such that noise from other confounding factors obscures any dependence of performance on tree-likeness.
@@ 5.1 Measuring Textural Separability
We will define the textural contrast or separability between some object mask and its surroundings by how easily their textures can be distinguished from one another. Motivated by findings that early layers of classification-pretrained CNNs primarily capture low-level features involving edges and textures [47], we characterize an image’s textures using the first convolutional layer of a ResNet-18 [18] pretrained on ImageNet [9]. Denoted by $f_{1}:\mathbb{R}^{C\times H\times W}\rightarrow\mathbb{R}^{C^{\prime}\times H^{\prime}\times W^{\prime}}$, this outputs a textural feature map corresponding to all $7\times 7$ windows in the image.
We then measure the textural separability of an object according to if a simple classifier $g$ can be trained to discriminate between (a) the activations for the pixels of the object foreground and (b) the activations right outside of the object boundary. We define this explicitly in Algorithm 3, where $\mathrm{Dilate}$ and $\mathrm{Disk}$ refer to scikit-image.morphology functions [43]. We use simple logistic regression for $g$ with inverse regularization parameter $C=2$ [32]. We also evaluate using other hyperparameter settings, as well as a random forest classifier for $g$ instead, in Appendix B.2, where we found similar results.
@@
7:  return  Training classification accuracy of $g$.
@@ 5.2.1 Style Transfer for Controlled Textural Contrast
To isolate the effects of textural separability from object shape, we first conduct controlled experiments using neural style transfer (NST) [14], before evaluating on real data in the next section. This allows us to systematically adjust an object’s textural contrast with its background while keeping other factors controlled.
We begin with all objects sampled from VOC2012 [12] that occupy 5–25% of their image area, to ensure sufficient background context (484 objects total). Using a Stable Diffusion-based inpainting model (App. C.2), we remove each object and reconstruct its background, creating pairs of separate object-only and background-only images. This setup allows precise modifications before reinserting objects into their original scenes. For each object, we then create three types of composite images with varying degrees of changes to the object’s shape (without modifying the background):
@@
After creating these image variants, we apply NST (similar to [15]) using textures randomly sampled from Colored Brodatz [1] with the Gatys et al. NST model [14] (details in App. C.2). An example of this entire procedure is in Fig. 6.
@@
In Fig. 8 right, we analyze how SAM’s performance varies with NST intensity across the three object transformation types. First, as textural separability decreases (higher NST intensity), SAM’s performance declines, confirming the impact of textural contrast on segmentation accuracy. Second, we observe clear performance gaps between transformation types: SAM performs best on altered objects, which undergo both shape and texture distortion, followed closely by mixed objects, and then performs worst on controlled objects. This suggests that additional textural contrast—introduced by shape, boundary, and texture transformations—enhances segmentation performance, further reinforcing the role of textural cues in SFM behavior.
@@ 5.2.2 Experiments on Real Data
We will now evaluate datasets of real images; iShape (Sec. 4.3.2) and Plittersdorf [17], both of which were used in the original SAM paper [26], and possess objects with a wide range of textural separability from their surroundings. Plittersdorf consists of camera trap video frames of wild deer recorded in a wildlife park. These frames sometimes have low-contrast objects due to frequent low-light conditions, making it a useful dataset to analyze. Example images and objects from both datasets are shown in App. A.4.
In Fig. 9, we show how the textural separability (Algorithm 3) of these objects relates to SAM’s performance on them, with correlation results shown for all three SFMs in Table 3. Overall, across all datasets, we see a fairly strong correlation between textural separability and segmentation IoU (average corr. of $\tau=0.47$ and $\rho=0.62$), especially considering the variety of objects and backgrounds, with objects with low separability typically being noticeably harder to segment. In these cases, the SFMs were confused by objects close to the object of interest which also have similar textures.
@@
antenna branch fence hanger log wire Plittersdorf
Model $\tau$ $\rho$ $\tau$ $\rho$ $\tau$ $\rho$ $\tau$ $\rho$ $\tau$ $\rho$ $\tau$ $\rho$ $\tau$ $\rho$
SAM, ViT-H [26] 0.44 0.59 0.50 0.68 0.65 0.85 0.63 0.83 0.33 0.48 0.49 0.67 0.26 0.38
SAM, ViT-B [26] 0.46 0.61 0.56 0.75 0.67 0.86 0.65 0.84 0.36 0.52 0.55 0.73 0.36 0.50
SAM 2, ViT-L [36] 0.46 0.64 0.56 0.76 0.64 0.83 0.65 0.84 0.25 0.36 0.54 0.71 0.20 0.29
SAM 2, ViT-B+ [36] 0.44 0.61 0.48 0.67 0.69 0.88 0.64 0.83 0.20 0.28 0.50 0.67 0.12 0.17
HQ-SAM, ViT-H [24] 0.39 0.54 0.50 0.68 0.69 0.88 0.62 0.80 0.20 0.28 0.47 0.63 0.20 0.28
HQ-SAM, ViT-B [24] 0.43 0.60 0.50 0.68 0.68 0.86 0.62 0.81 0.23 0.33 0.49 0.66 0.21 0.31
Average 0.44 0.60 0.52 0.70 0.67 0.86 0.64 0.83 0.26 0.38 0.51 0.68 0.23 0.32
Table 3: Rank correlation between SFM IoU and object textural separability, on iShape and Plittersdorf.
Figure 9: Segmentation IoU vs. object textural separability, for iShape (left) shown for SAM, and Plittersdorf (right) for all SFMs.
@@ 6 Robustness to Fine-tuning
Despite being designed for thin, branching structures, HQ-SAM exhibits performance drops on high tree-likeness objects, similar to SAM and SAM 2 (Table 2, Fig. 5). This suggests that the difficulty SFMs face with such objects persists even after adaptation, a hypothesis we explore here.
To assess whether fine-tuning could improve SAM’s segmentation of high tree-likeness objects, we selected some $k$ examples from DIS-train using different ranking strategies: Oracle (direct IoU-based selection), Rand (random selection), DoGD / CPR (selecting examples based on high tree-likeness), and Sum (a combination of DoGD and CPR). We fine-tuned only SAM’s mask decoder on these, keeping the image encoder frozen to preserve its generalist capabilities (as in [24]); full details are in App. C.3.
@@ 7 Discussion
Understanding SFM Failure Modes. We find that SFMs struggle primarily with over-segmentation rather than under-segmentation in high tree-likeness cases (Sec. 4.3). The models often generate large false-positive regions within densely packed, thin structures (e.g., Figs. 3, right; 1, left). This cannot be explained by resolution limitations alone, as sparsely packed thin structures exhibit reduced over-segmentation (Fig. 3, center rightmost). Instead, we hypothesize that SFMs misinterpret dense, irregular patterns as texture, confusing these structures with the appearance of a more uniformly shaped object.
In the cases of both tree-like and low-textural contrast objects, SFMs struggle to delineate object boundaries when faced with misleading textural cues, be it either a dense, irregular shape that appears to instead be a texture (the former case), or an object texture that is similar to its surroundings (the latter). This suggests that SAM’s training set (SA-1B) contained few such objects, as tree-like structures and low-contrast objects are uncommon in typical photographic datasets [26] (for the former, see e.g., the typically low object concavity of SA-1B in Fig. 6, right in [26]). Since HQ-SAM and SAM 2 were derived from SAM/its training set, their similar behavior is unsurprising.
On HQ-SAM and Fine-tuning. Despite being designed and adapted for thin, intricate structures [24], HQ-SAM still exhibits performance drops correlated with tree-likeness (Tables 1, 2), albeit with slight improvements over SAM and SAM 2 (Fig. 5, left). This may stem from HQ-SAM’s adapter-based approach, which leaves SAM’s image encoder and mask decoder frozen, preserving SAM’s limitations. Our fine-tuning experiments (Sec. 6) further suggest that this issue is not easily mitigated, indicating that these structures pose a fundamental segmentation challenge.
@@
Conclusion. In this paper, we quantitatively modeled how SFM segmentation performance relates to tree-likeness and textural separability. Our findings reveal a consistent correlation between these object characteristics and segmentation accuracy, highlighting the need for further research to better understand and mitigate these failure modes—especially since even targeted fine-tuning fails to fully resolve them.
@@ A.1 Retinal Blood Vessel and Road Satellite Images
The retinal blood vessel and road satellite image object masks which we use to generate the synthetic images for the experiments in Sec. 4.3.1, are from the Retina Blood Vessel [22] and the Road Extraction Challenge data of DeepGlobe18 [8]. For our retinal vessel mask set, we use the vessel masks from the training set of 80 image/mask pairs, and for our road mask set, we randomly sample the same number of masks from the split-merged full dataset, in order to ensure that these two object types appear equally in the synthetic dataset of Sec. 4.3.1.

### v3

@@ 4.3.2 Experiments on Real Data
DIS MOSE antenna branch fence hanger log wire
Model $\tau$ $\rho$ $\tau$ $\rho$ $\tau$ $\rho$ $\tau$ $\rho$ $\tau$ $\rho$ $\tau$ $\rho$ $\tau$ $\rho$ $\tau$ $\rho$
CPR:
SAM-H -0.59 -0.76 -0.80 -0.95 -0.13 -0.21 -0.61 -0.81 -0.61 -0.81 -0.25 -0.37 -0.60 -0.77 -0.06 -0.10
SAM-B -0.63 -0.81 -0.76 -0.93 -0.18 -0.27 -0.63 -0.82 -0.63 -0.83 -0.44 -0.62 -0.62 -0.80 -0.12 -0.17
SAM 2-L -0.41 -0.57 -0.84 -0.96 -0.30 -0.43 -0.64 -0.84 -0.61 -0.81 -0.48 -0.66 -0.67 -0.85 -0.14 -0.21
SAM 2-B+ -0.38 -0.52 -0.81 -0.95 -0.16 -0.23 -0.52 -0.71 -0.57 -0.78 -0.42 -0.59 -0.60 -0.77 -0.12 -0.18
HQ-SAM-H -0.52 -0.69 -0.79 -0.94 -0.23 -0.33 -0.77 -0.93 -0.79 -0.95 -0.51 -0.70 -0.69 -0.85 -0.22 -0.33
HQ-SAM-B -0.62 -0.81 -0.76 -0.92 -0.25 -0.34 -0.75 -0.92 -0.83 -0.96 -0.50 -0.69 -0.71 -0.87 -0.24 -0.35
Average -0.53 -0.69 -0.79 -0.94 -0.21 -0.30 -0.65 -0.84 -0.67 -0.86 -0.43 -0.61 -0.65 -0.82 -0.15 -0.22
DoGD:
SAM-H 0.58 0.77 0.43 0.60 0.23 0.30 0.52 0.70 0.46 0.65 0.52 0.72 0.30 0.46 0.08 0.11
SAM-B 0.45 0.64 0.27 0.39 0.12 0.14 0.51 0.69 0.45 0.64 0.50 0.70 0.13 0.22 0.06 0.08
SAM 2-L 0.44 0.60 0.52 0.70 0.42 0.59 0.48 0.65 0.49 0.68 -0.05 -0.08 0.56 0.75 -0.03 -0.05
SAM 2-B+ 0.26 0.39 0.47 0.66 0.40 0.56 0.29 0.40 0.30 0.44 -0.06 -0.10 0.42 0.58 -0.08 -0.12
HQ-SAM-H 0.48 0.66 0.41 0.58 0.53 0.72 0.58 0.72 0.48 0.67 -0.31 -0.45 0.57 0.75 -0.13 -0.19
HQ-SAM-B 0.50 0.69 0.33 0.48 0.67 0.85 0.59 0.73 0.57 0.77 -0.25 -0.34 0.53 0.71 -0.04 -0.06
Average 0.45 0.63 0.41 0.57 0.40 0.53 0.50 0.65 0.46 0.64 0.06 0.08 0.42 0.58 -0.02 -0.04
Figure 5: Left: Segmentation IoU vs. object tree-likeness (CPR, top; DoGD, bottom) on DIS (left) for all three SFMs and iShape (right) shown for SAM-H. Right: rank correlations between IoU and tree-likeness for all SFMs.
We see that despite the many potential confounding factors in real data, there is still a clear correlation between object tree-likeness as measured by the proposed metrics and SFM segmentation performance. In particular, we see average absolute correlations of $|\tau|=0.62$ and $|\rho|=0.79$ for CPR (particularly high for MOSE) and $|\tau|=0.38$ and $|\rho|=0.53$ for DoGD—excluding the antenna and wire objects of iShape which had outlying correlations likely because those two object classes cover only small ranges of tree-likeness as opposed to the other iShape classes (Fig. 5, left), such that noise from other confounding factors obscures any dependence of performance on tree-likeness.
@@ 5.1 Measuring Textural Separability
We will define the textural contrast or separability between some object mask and its surroundings by how easily their textures can be distinguished from one another. Motivated by findings that early layers of classification-pretrained CNNs primarily capture low-level features involving edges and textures [zeiler2014visualizing], we characterize an image’s textures using the first convolutional layer of a ResNet-18 [he2016resnet] pretrained on ImageNet [deng2009imagenet]. Denoted by $f_{1}:\mathbb{R}^{C\times H\times W}\rightarrow\mathbb{R}^{C^{\prime}\times H^{\prime}\times W^{\prime}}$, this outputs a textural feature map corresponding to all $7\times 7$ windows in the image.
We then measure the textural separability of an object according to if a simple classifier $g$ can be trained to discriminate between (a) the activations for the pixels of the object foreground and (b) the activations right outside of the object boundary. We define this in Algorithm 3, where $\mathrm{Dilate}$ and $\mathrm{Disk}$ refer to scikit-image.morphology functions [scikit-image]. We use simple logistic regression for $g$ with regularization parameter $C=2$ [scikit-learn]. We also evaluate other hyperparameter settings, as well as a random forest classifier for $g$, in App. B.2, where we found similar results.
@@
7:  Return Test classification accuracy of $g$.
@@ 5.2.1 Style Transfer for Controlled Textural Contrast
To isolate the effects of textural separability from object shape, we first conduct controlled experiments using neural style transfer (NST) [gatys2016image], before evaluating on real data in the next section. This allows us to systematically adjust an object’s textural contrast with its background while keeping other factors controlled.
We begin with all objects sampled from VOC2012 [pascal-voc-2012] that occupy 5–25% of their image area, to ensure sufficient background context (484 objects total). Using a Stable Diffusion-based inpainting model (App. C.2), we remove each object and reconstruct its background, creating pairs of separate object-only and background-only images. This setup allows precise modifications before reinserting objects into their original scenes. We then create three types of composite images with varying degrees of changes to object shape for each object (without modifying the background):
@@
After creating these image variants, we apply NST (similar to [geirhosimagenet2019]) using textures randomly sampled from Colored Brodatz [abdelmounaime2013newMBT] with the Gatys et al. NST model [gatys2016image] (details in App. C.2). An example of this entire procedure is in Fig. 6.
@@
In Fig. 8 right, we analyze how SAM’s performance varies with NST intensity across the three object transformation types (we found similar results for the other SFMs, shown in App. D.5). First, as textural separability decreases (higher NST intensity), SAM’s performance declines, confirming the impact of textural contrast on segmentation accuracy. Second, we observe clear performance gaps between transformation types: SAM performs best on altered objects, which undergo both shape and texture distortion, followed closely by mixed objects, and then performs worst on controlled objects. This suggests that additional textural contrast–introduced by shape, boundary, and texture transformations–enhances segmentation performance, further reinforcing the role of textural cues in SFM behavior.
@@ 5.2.2 Experiments on Real Data
We will now evaluate datasets of real images: iShape, MOSE (Sec. 4.3.2), and Plittersdorf [haucke2022socrates]—of which iShape and Plittersdorf were used in the original SAM paper [sam]—which all possess objects with a wide range of textural separability from their surroundings. Plittersdorf consists of camera trap video frames of wild deer recorded in a wildlife park. These frames sometimes have low-contrast objects due to frequent low-light conditions, making it a useful dataset to analyze. Example images and objects from both datasets are shown in App. A.4.
In Fig. 9 left, we show how the textural separability (Algorithm 3) of these objects relates to SAM’s performance on them, with correlation results shown for all three SFMs to the right of the plots (MOSE plots shown in App. D.7). Overall, across all datasets, we see a fairly strong correlation between textural separability and segmentation IoU (average corr. of $\tau=0.49$ and $\rho=0.65$), especially considering the variety of objects and backgrounds, with objects with low separability typically being noticeably harder to segment. This relationship was especially strong for the MOSE dataset, with average $\tau=0.65$ and $\rho=0.82$. In these cases, the SFMs were confused by objects close to the object of interest which also have similar textures.
Refer to captionRefer to caption
@@
antenna branch fence hanger log wire Plitters. MOSE
Model $\tau$ $\rho$ $\tau$ $\rho$ $\tau$ $\rho$ $\tau$ $\rho$ $\tau$ $\rho$ $\tau$ $\rho$ $\tau$ $\rho$ $\tau$ $\rho$
SAM-H 0.44 0.59 0.50 0.68 0.65 0.85 0.63 0.83 0.33 0.48 0.49 0.67 0.26 0.38 0.65 0.83
SAM-B 0.46 0.61 0.56 0.75 0.67 0.86 0.65 0.84 0.36 0.52 0.55 0.73 0.36 0.50 0.74 0.90
SAM 2-L 0.46 0.64 0.56 0.76 0.64 0.83 0.65 0.84 0.25 0.36 0.54 0.71 0.20 0.29 0.58 0.76
SAM 2-B+ 0.44 0.61 0.48 0.67 0.69 0.88 0.64 0.83 0.20 0.28 0.50 0.67 0.12 0.17 0.58 0.75
HQ-SAM-H 0.39 0.54 0.50 0.68 0.69 0.88 0.62 0.80 0.20 0.28 0.47 0.63 0.20 0.28 0.64 0.82
HQ-SAM-B 0.43 0.60 0.50 0.68 0.68 0.86 0.62 0.81 0.23 0.33 0.49 0.66 0.21 0.31 0.72 0.88
Average 0.44 0.60 0.52 0.70 0.67 0.86 0.64 0.83 0.26 0.38 0.51 0.68 0.23 0.32 0.65 0.82
Figure 9: Left: Segmentation IoU vs. object textural separability, for iShape (left) shown for SAM-H, and Plittersdorf (right) for all SFMs. Right: Rank correlations between SFM IoU and object textural separability, on iShape, Plittersdorf and MOSE for all SFMs.
@@ 6 Robustness to Fine-tuning
Despite being designed for thin, branching structures, HQ-SAM exhibits performance drops on high tree-likeness objects, similar to SAM and SAM 2 (Fig. 5). This suggests that the difficulty SFMs face with such objects persists even after adaptation, a hypothesis we explore here.
To assess whether fine-tuning could improve SAM’s segmentation of high tree-likeness objects, we selected some $k$ examples from DIS-train using different ranking strategies: Oracle (direct IoU-based selection), Rand (random selection), DoGD / CPR (selecting examples based on high tree-likeness), and Sum (a combination of DoGD and CPR). We fine-tuned only SAM’s mask decoder on these, keeping the image encoder frozen to preserve its generalist capabilities (as in [hqsam]); full details are in App. C.3.
@@ 7 Discussion
Understanding SFM Failure Modes. We find that SFMs struggle primarily with over-segmentation rather than under-segmentation in high tree-likeness cases (Sec. 4.3), generating large false-positive regions within densely packed, thin structures (e.g., Figs. 3, right; 1, left). This cannot be explained by resolution limitations alone, as sparse thin structures exhibit reduced over-segmentation (Fig. 3, center rightmost). Instead, we hypothesize that SFMs misinterpret dense, irregular patterns as texture, confusing these structures with the appearance of a more uniformly-shaped object.
For both tree-like and low-contrast objects, SFMs struggle to delineate object boundaries when faced with misleading textural cues, be it either a dense, irregular shape that appears to instead be a texture (the former case), or an object texture similar to its surroundings (the latter). This suggests that SAM’s training set (SA-1B) contained few such objects, as tree-like structures and low-contrast objects are uncommon in typical photographic datasets [sam] (for the former, see e.g., the typically low object concavity of SA-1B in Fig. 6, right in [sam]). Since HQ-SAM and SAM 2 were derived from SAM, their similar behavior is unsurprising.
Comparison to Non-SFM Models. Additional experiments with models trained from scratch (IS-Net, iShapeInst, Mask R-CNN) and evaluated on the DIS, iShape and Plittersdorf datasets, respectively (App. D.6), revealed that their performance did not sharply degrade on objects with high tree-likeness or low textural separability. In contrast to SFMs, such objects were not consistent “failure modes” for these models, suggesting that these difficulties are specific to SFMs rather than being inherently hard cases.
On HQ-SAM and Fine-tuning. Despite being designed and adapted for thin, intricate structures [hqsam], HQ-SAM still exhibits performance drops correlated with tree-likeness (Figs. 4, 5), albeit with slight improvements over SAM and SAM 2 (Fig. 5, left). This may stem from HQ-SAM’s adapter-based approach, which leaves SAM’s image encoder and mask decoder frozen, preserving SAM’s limitations. Our fine-tuning experiments (Sec. 6) further suggest that this issue is not easily mitigated, indicating that these structures pose a fundamental segmentation challenge.
@@ A.1 Retinal Blood Vessel and Road Satellite Images
The retinal blood vessel and road satellite image object masks which we use to generate the synthetic images for the experiments in Sec. 4.3.1, are from the Retina Blood Vessel [noauthor_retina_nodate] and the Road Extraction Challenge data of DeepGlobe18 [DeepGlobe18]. For our retinal vessel mask set, we use the vessel masks from the training set of 80 image/mask pairs, and for our road mask set, we randomly sample the same number of masks from the split-merged full dataset, in order to ensure that these two object types appear equally in the synthetic dataset of Sec. 4.3.1.